OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least 3 sentences

OpenStudy (anonymous):

will medal and fan

OpenStudy (anonymous):

The differences are in the y-intercept, the 2x+3y-1200 is transformed to y= -2/3 x+400. Also 2x+3y=1500 and the intercept is 500 and the slope being -2/3.

OpenStudy (anonymous):

thanks so much man!